OpenStudy (anonymous):

After Phillip's birthday party, there were several soft-drink bottles with some soda left inside. There were 5 bottles each with 4 ounces left, 2 bottles each with 6 ounces left, and 3 bottles each with 8 ounces left. What is the average number of ounces left per bottle? 5 ounces 6 ounces 5.6 ounces 4.1 ounces

OpenStudy (anonymous):

@texaschic101 help quick

OpenStudy (anonymous):

add all of the ounces together and divide by the number of bottles (56÷10)

OpenStudy (anonymous):

what do you get?

OpenStudy (texaschic101):

5(4) + 2(6) + 3(8) = 20 + 12 + 24 = 56 oz 5 bottles + 2 bottles + 3 bottles = 10 bottles 56/10 = 5.6 oz

OpenStudy (anonymous):

5.6

OpenStudy (texaschic101):

correct :)

OpenStudy (anonymous):

right. which choice is that?

OpenStudy (anonymous):

option 3 or also known as C thanks

OpenStudy (anonymous):

Correct :) <3 no problem

OpenStudy (texaschic101):

nicely done Rosie :)

OpenStudy (anonymous):

thanks @texaschic101 :)